Q. Find the domain of the function y=√(2-√x) .
1
Expert's answer
2019-07-19T09:25:10-0400

Solution:


y="\sqrt{\smash[b]{2-\sqrt{\smash[b]{x}}}}"

Since we have square root ,So the value under this cannot be negative.


It means 2-"\sqrt{\smash[b]{x}}" 0  ------------- condition (1)


However at the same time we need to take care that "\sqrt{\smash[b]{x}}"   0 ------------- condition (2)


So considering both the condition we can say x ≥ 0 and x ≤ 4


Let x=0

y = "\sqrt{\smash[b]{2-\sqrt{\smash[b]{0}}}}" ⟹ valid number under square root


Let x=1

y = "\sqrt{\smash[b]{2-\sqrt{\smash[b]{1}}}}" ⟹ valid number under square root


Let x=2

y = "\sqrt{\smash[b]{2-\sqrt{\smash[b]{2}}}}" ⟹ valid number under square root


Let x=3

y = "\sqrt{\smash[b]{2-\sqrt{\smash[b]{3}}}}" ⟹ valid number under square root


Let x= 4

y = "\sqrt{\smash[b]{2-\sqrt{\smash[b]{4}}}}" ⟹ valid number under square root


Now Let x=5

y = "\sqrt{\smash[b]{2-\sqrt{\smash[b]{5}}}}" ⟹ This gives a negative number under square root


From here(x ≥ 5) each and every value of x will give a negative value under square root


So Domain of function is [0,4]